# CAPE: Covariate-Adjusted Pre-Training for Epidemic Time Series Forecasting

### v3

@@ 3.1 Causal Analysis for Epidemic Forecasting
$p(Y|do(X))=\sum\nolimits_{Z}p(Y|X,Z=z)p(Z=z).$ (2)
However, obtaining detailed environmental information, or ${\bf e}_{k}$, can be challenging due to variability in data availability and quality. To address this, we resort to a data-driven approach that treats ${\bf e}_{k}$ as learnable parameters and thus allows us to dynamically infer the environmental distribution directly from the observed data. Specifically, we implement an environment estimator $q_{\phi}(Z|X)$ that infers the probability of environment states based on historical inputs together with the latent representations of each state. Then, a variational lower bound is derived (Wu et al., 2024) (see Appendix A.3):
$\log{p_{\Theta}(\hat{Y}|do(X))}\geq$ $\mathbb{E}_{q_{\phi}(Z|X)}\left[\log p_{\Theta}(\hat{Y}|X,Z)\right]$ (3)
$-\operatorname{KL}\left(q_{\phi}(Z|X)\parallel p(Z)\right),$
where the first term maximizes the model’s predictive power and the second term regularizes the environment estimator to output a distribution close to the prior distribution $p(Z)$.
3.2 Model Instantiation
Figure 3: (a) CAPE encoder and environment estimator with latent representations; (b) Hierarchical environment contrasting for temporal and environment representations; (c) Random masking and reconstruction with environment estimation to capture universal patterns; (d) EM algorithm to iteratively optimize model parameters and environment representations.
To instantiate and train a model that performs the covariate adjustment, we need to model the environment estimator $q_{\phi}(Z|X)$ and the predictor $p_{\Theta}(\hat{Y}|X,Z)$.
Latent Environment Estimator $q_{\phi}(Z|X)$. We approximate the distribution of environmental contexts $p(Z)$ using a latent environment estimator $q_{\phi}(Z|X)$. Epidemic infection trajectories are profoundly shaped by dynamic environmental influences—such as the implementation and lifting of non-pharmaceutical interventions (NPIs), seasonal drivers affecting pathogen viability (e.g., influenza seasonality), or shifts in population behavior—which can vary significantly over different periods of an outbreak. Modeling the impact of these potentially complex and time-varying influences across an entire long infection trajectory can be inefficient and may obscure shorter-term but critical environmental shifts. Therefore, we apply patching (Nie et al., 2022) to manage the granularity in our environment estimation. This approach allows the model to identify distinct environmental ”regimes” or periods within an epidemic wave that might correspond to different dominant external factors (e.g., a period of strict lockdown versus a period of reopening). Such granularity helps prevent overly coarse or overly fine-grained estimations that could obscure epidemiologically relevant shifts in transmission dynamics, such as changes in the effective reproduction number or the shape of the epidemic curve. We divide the historical infection input $\mathbf{x}$ (e.g., weekly case counts) into $C$ non-overlapping temporal patches, $\mathbf{x}=[\mathbf{x}_{1},\dots,\mathbf{x}_{C}]$, where $\mathbf{x}_{c}\in\mathbb{R}^{T/C}$. A self-attention layer $f_{\text{enc}}$ then captures temporal dependencies between these patches, producing contextualized representations of the local infection trend $\mathbf{h}^{(l)}_{c}=f_{\text{enc}}(\mathbf{x}^{(l)}_{c})$ for each patch at layer $l$. Subsequently, the environment state and the patch representation are projected to the same dimension and we calculate the cross-attention score using a softmax function as the environment probability:
where $\pi_{k,c}^{(l)}=q_{\phi}(Z=z_{k}|X=x_{c})$ is the output probability of the environment $z_{k}$ for the $c$-th patch, and $\mathbf{W}_{k}^{(l)}$ and $\mathbf{W}_{s}^{(l)}$ are learnable linear layers. This mechanism allows the model to not only leverage the observed infection pattern within the patch $\mathbf{h}_{c}^{(l)}$ but also to consider a set of learned environmental influences $\textbf{e}_{k}$. The $\textbf{e}_{k}$ are optimized to capture distinct types of environmental conditions, expressed as a vector of channel-wise weights, relevant to disease spread across many historical epidemics encountered during pre-training. This design is crucial as it enables the model to potentially infer the influence of combinations or intensities of environmental factors not explicitly seen together in the training data for a specific downstream task, thereby enhancing its ability to generalize to novel epidemic scenarios or evolving outbreaks.
Epidemic Predictor $p_{\Theta}(\hat{Y}|X,Z)$. Unlike previous studies, which do not explicitly model environment states, we incorporate these states directly into the patched infection trend using their latent representations $\mathbf{e}_{k}$. Specifically, we model the predictor $p_{\Theta}(\hat{Y}|X,Z)$ in Eq (2) by employing a weighted sum over the combined representations of each environment and the infection trend using Hadamard product, i.e., $f_{\text{enc}}(\mathbf{x}_{c}^{(l)})\odot\mathbf{e}_{k}$. Finally, we apply a feed-forward layer is applied. This step gives us the adjusted environment-invariant representation for $X_{ca}$ in Figure 2, and we conduct such adjustment in a layer-wise manner, which can be expressed as:
$\small\mathbf{x}^{(l+1)}_{c}=\sigma\left(\mathbf{W}_{f}^{(l)}\sum\nolimits_{k=1}^{K}\pi_{k,c}^{(l)}\left[f_{enc}({\mathbf{x}_{c}^{(l)}})\odot\mathbf{e}_{k}\right]\right),$ (5)
where $\sigma$ represents the activation function and $\mathbf{W}_{f}^{(l)}$ denotes the learnable parameters of the feedforward layer. Assuming $L$ layers are stacked, we acquire the final representation $\mathbf{x}^{(L)}=[\mathbf{x}^{(L)}_{1},\mathbf{x}^{(L)}_{2},\ldots\mathbf{x}^{(L)}_{C}]=g_{\theta}({\bf x})\in\mathbb{R}^{C\cdot d}$ and apply a task-specific head to predict the target variable $\mathbf{y}=h_{\psi}(\mathbf{x}^{(L)})$, where $h_{\psi}$ is a linear transformation.
3.3 Pre-training Objectives for Epidemic Forecasting
CAPE captures diverse epidemic time series dynamics through self-supervised learning tasks that identify universal patterns in the pre-training dataset. While previous studies neglected the confounding effects of environmental factors on input-label pairs in $\mathcal{T}_{\text{pre}}$, CAPE seamlessly integrates environment estimation into the self-supervised framework.
Random Masking with Environment Estimation. To learn from large unlabeled epidemic time series, we adopt a masked modeling task (Kamarthi & Prakash, 2023; Goswami et al., 2024), masking 30% of input patches (Figure 3(c)). As shown in Figure 2, input $X$ is influenced by the environment $Z$, so reconstructing patches requires modeling both temporal and environmental dependencies. Unlike previous methods that ignore this, we use backdoor adjustment via an environment estimator $q_{\phi}(Z|X)$ to obtain environment-invariant representations. During pre-training, the input $\mathbf{x}$ is masked to produce pairs $(\tilde{\mathbf{x}},\mathbf{x})$, with the reconstruction $\hat{\mathbf{x}}=h_{\psi}(g_{\theta}(\tilde{\mathbf{x}}))$ trained using Mean Squared Error: $\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})=\text{MSE}(\hat{\mathbf{x}},\mathbf{x})$.
Hierarchical Environment Contrasting for Robust Epidemiological Contextualization. During optimization, inconsistent environment representations can occur if the model, when processing two infection trajectories (${\bf x}$ and ${\bf x}^{\prime}$) that share an identical overlapping time period, estimates different environmental contexts for that common period based on the distinct temporal context. To address the challenge and improve consistency in representing environments for adjacent epidemic time series segments, we propose Hierarchical Environment Contrasting. Specifically, We define an aggregated environment representation as the confluence of environmental factors, $\hat{\mathbf{e}}^{(l)}_{c}=\sum_{k=1}^{K}\mathbf{e}_{k}\pi_{k,c}^{(l)}$, for each patch $c$, allowing it to reflect a nuanced, weighted combination of $K$ learned latent environments ($\mathbf{e}_{k}$) rather than a single discrete state. Our contrastive feature, $\hat{\mathbf{E}}_{j,c}^{(l)}=\sigma(\mathbf{W}_{f}^{(l)}(\hat{\mathbf{e}}^{(l)}_{c}\odot\mathbf{h}^{(l)}_{c}))$, is designed to capture this essential interaction. It explicitly combines the aggregated environment representation ($\hat{\mathbf{e}}^{(l)}_{c}$) with the local infection trend representation ($\mathbf{h}^{(l)}_{c}$) via a Hadamard product ($\odot$), enabling the model to learn how specific environmental profiles alter observed infection trends (e.g., by changing growth rates). The overarching goal, realized through the contrastive loss (Equation 3.3), is to ensure these learned features $\hat{\mathbf{E}}_{j,c}^{(l)}$ are consistent across genuinely comparable epidemiological contexts and distinct for dissimilar underlying situations, even if superficial case counts appear similar, fostering robust, discriminative representations of environmental influence. The patch-wise contrastive loss is:
$\mathcal{L}_{\text{CL}}(j,c)=-\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,c)}$
$\hskip 28.45274pt+\log\left(\sum\nolimits_{b\in B}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(b,c)}\right)+\mathbb{I}_{j\neq b}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(b,c)}\right)\right)$
$\hskip 28.45274pt+\log\left(\sum\nolimits_{t\in\Omega}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}^{\prime}}_{(j,t)}\right)+\mathbb{I}_{c\neq t}\exp\left(\hat{\mathbf{E}}_{(j,c)}\cdot\hat{\mathbf{E}}_{(j,t)}\right)\right).$
where $B$ is the batch, $\Omega$ is the set of overlapping patches, and $\mathbb{I}$ is the indicator function. For more details, please refer to Appendix A.4.
Given a batch $\mathbf{X}\in\mathbb{R}^{B\times T}$, our pre-training loss is: $\mathcal{L}_{\text{final}}=\sum\nolimits_{\mathbf{x}\in\mathbf{X}}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})+\alpha\,\mathcal{L}_{\text{CL}}(\hat{\mathbf{E}}^{(L)},\hat{\mathbf{E}}^{\prime(L)})$, where $L$ is the number of layers and $\alpha$ balances reconstruction and contrastive losses.
3.4 Optimization of the CAPE Framework
To maximize the variational lower bound in Eq. (3), we adopt the Expectation-Maximization (EM) algorithm to iteratively update latent environments and the epidemic predictor (see Appendix A.5 for the pseudo algorithm).
E-Step: Estimating Latent Environments. We estimate the environment states $Z$ and their distribution $p(Z)$ by maximizing the expected likelihood of $p(Y|Z)$. Fixing the epidemic predictor $p_{\Theta}(\hat{Y}|X,Z)$ and environment estimator $q_{\phi}(Z|X)$, we treat them as oracles such that $p_{\Theta}(\hat{Y}|X,Z)=p(Y|X,Z)$ and $q_{\phi}(Z|X)=p(Z)$. The environment representations $\mathbf{E}$ are updated by minimizing the expected reconstruction loss: $\mathbf{E}^{t+1}\leftarrow\arg\min\nolimits_{\mathbf{E}}\left[\mathbb{E}_{\mathbf{x}\sim p(X)}\mathcal{L}_{\text{recon}}(\mathbf{x},\hat{\mathbf{x}})\right]$. This yields the updated distribution $p^{t+1}(Z)=q^{t+1}_{\phi_{t}}(Z)$ and representations $\mathbf{E}^{t+1}$.
M-Step: Optimizing Epidemic Predictor. With $\mathbf{E}^{t+1}$ fixed, we provide the following theorem:
Theorem 3.1.
Assuming $q^{t+1}_{\phi_{t}}(Z)=p^{t+1}(Z)$ and an L2 norm is applied on $\phi$, the variational lower bound in Eq. (3) can be approximated as follows: $\mathbb{E}_{p(X)}\left[\mathbb{E}_{q^{t+1}_{\phi_{t}}(Z|X)}\left[\log p^{t+1}_{\Theta_{t+1}}(\hat{Y}|X,Z)\right]\right]-C$, which is equivalent to minimizing the expected reconstruction loss $\mathbb{E}_{\mathbf{x}\sim p(X)}[\mathcal{L}_{recon}(\mathbf{x},\hat{\mathbf{x}})]$
Further theoretical details are provided in Appendix A.3. Therefore, we optimize model parameters by minimizing the expected reconstruction loss and incorporating a contrastive loss to regularize the environment estimator: $\Theta_{t+1}\leftarrow\arg\min\nolimits_{\Theta}\left[\mathcal{L}_{\text{final}}(\mathbf{X},\hat{\mathbf{X}},\mathbf{E}^{t+1})\right]$.
@@ 3.4 Optimization of the CAPE Framework
Table 1: Univariate forecasting results with horizons ranging from 1 to 16 future steps. The lookback window length is set to 36 and all models are evaluated using MSE. Note that performance rankings are distinguished by color coding: Best, Second Best, Third Best. $\Delta(\%)$ stands for the relative improvement of CAPE over the baselines in terms of average MSE over all horizons.
Dataset Horizon Statistical Model RNN-Based MLP-Based Transformer-Based CAPE
ARIMA LSTM GRU Dlinear Non-Pre-trained Pre-trained
Informer Autoformer Fedformer PEM MOMENT PatchTST
ILI USA 1 0.138 0.338 0.259 0.220 0.175 0.457 0.368 0.179 0.269 0.195 0.155
2 0.203 0.377 0.301 0.247 0.370 0.710 0.380 0.226 0.321 0.264 0.200
4 0.354 0.458 0.386 0.376 0.517 0.670 0.433 0.304 0.397 0.385 0.270
8 0.701 0.579 0.529 0.506 0.597 0.842 0.570 0.538 0.510 0.535 0.404
16 1.121 0.691 0.626 0.617 0.812 0.835 0.701 0.570 0.610 0.485 0.516
Avg 0.503 0.489 0.420 0.393 0.494 0.703 0.490 0.363 0.421 0.373 0.309
$\Delta$ (%) 38.57% 36.81% 26.43% 21.37% 37.45% 56.05% 36.94% 14.88% 26.60% 17.16% -
ILI Japan 1 0.358 1.426 1.213 1.016 0.405 0.515 0.525 0.470 0.325 0.413 0.290
2 0.772 1.635 1.458 1.294 0.666 0.855 1.151 0.755 0.586 0.698 0.535
4 1.720 1.975 1.870 1.758 1.234 1.150 1.455 1.207 1.082 1.147 0.944
8 2.981 2.373 2.365 2.285 1.688 1.866 2.012 1.810 1.706 1.708 1.650
16 2.572 2.023 2.010 2.007 1.551 2.654 4.027 1.766 2.054 1.688 1.911
Avg 1.680 1.886 1.783 1.672 1.109 1.408 1.834 1.202 1.151 1.131 1.066
$\Delta$ (%) 36.55% 43.48% 40.21% 36.24% 3.88% 24.29% 41.88% 11.31% 7.38% 5.74% -
Measles 1 0.071 0.182 0.143 0.133 0.066 0.203 0.321 0.085 0.113 0.094 0.083
2 0.120 0.223 0.176 0.184 0.153 0.257 0.817 0.128 0.138 0.127 0.112
4 0.225 0.310 0.258 0.296 0.288 0.331 0.226 0.213 0.186 0.205 0.161
8 0.483 0.567 0.471 0.512 0.501 0.671 0.403 0.417 0.351 0.377 0.310
16 1.052 1.110 1.013 1.088 0.904 1.115 0.754 0.806 0.818 0.722 0.752
Avg 0.390 0.478 0.412 0.443 0.382 0.515 0.504 0.330 0.321 0.305 0.269
$\Delta$ (%) 31.03% 43.72% 34.71% 39.28% 29.58% 47.77% 46.63% 18.49% 16.20% 11.80% -
Dengue 1 0.244 0.250 0.261 0.224 0.255 0.525 0.521 0.225 0.420 0.240 0.223
2 0.373 0.343 0.343 0.316 0.450 0.807 0.670 0.314 0.579 0.334 0.302
4 0.696 0.564 0.579 0.560 0.798 0.957 0.766 0.571 0.661 0.586 0.561
8 1.732 1.168 1.183 1.256 1.239 1.684 1.539 1.223 1.308 1.292 1.046
16 4.082 3.876 3.315 3.109 2.659 3.364 2.934 3.376 2.532 2.537 2.509
Avg 1.426 1.240 1.136 1.093 1.080 1.467 1.286 1.142 1.100 1.000 0.892
$\Delta$ (%) 37.45% 28.06% 21.48% 18.39% 17.41% 39.20% 30.64% 21.89% 18.91% 10.80% -
Covid 1 33.780 22.592 22.009 23.811 34.161 42.049 28.130 25.088 32.376 23.645 21.548
2 33.193 23.460 22.542 24.809 24.883 30.631 28.059 23.123 35.418 25.047 22.224
4 32.482 24.729 24.816 26.345 31.328 41.029 29.432 23.889 36.251 24.224 22.476
8 36.573 31.019 33.934 33.081 35.964 55.812 41.791 31.217 40.429 31.548 28.403
16 42.910 43.820 41.432 47.561 50.244 47.993 69.976 51.265 52.590 43.309 40.555
Avg 35.787 29.124 28.947 31.121 35.316 43.503 39.478 30.917 39.413 29.555 26.559
$\Delta$ (%) 25.79% 8.81% 8.25% 14.66% 24.80% 38.95% 32.72% 14.10% 32.61% 10.14% -

### v4

@@ 3.3 Epidemic-aware Regularization
[…]
6:Calculate proxy bounds: $\widehat{R}^{\mathrm{lo}}_{0}\leftarrow\frac{\sigma_{\min}(\mathbf{F})}{\sigma_{\max}(\mathbf{V})}$, $\widehat{R}^{\mathrm{hi}}_{0}\leftarrow\frac{\sigma_{\max}(\mathbf{F})}{\sigma_{\min}(\mathbf{V})}$.
7:Calibrate estimates $(\widehat{R}^{\mathrm{lo}}_{0},\widehat{R}^{\mathrm{hi}}_{0})\leftarrow\mathrm{Calib}_{\theta}(\widehat{R}^{\mathrm{lo}}_{0},\widehat{R}^{\mathrm{hi}}_{0})$ and compute the loss:
$\mathcal{L}_{R_{0}}\leftarrow\max\{0,\,R^{\mathrm{lo}}_{0}(d)-\widehat{R}^{\mathrm{hi}}_{0}\}+\max\{0,\,\widehat{R}^{\mathrm{lo}}_{0}-R^{\mathrm{hi}}_{0}(d)\}$
8:return $(\widehat{R}^{\mathrm{lo}}_{0},\widehat{R}^{\mathrm{hi}}_{0},\mathcal{L}_{R_{0}})$.
Compartment Alignment Loss. Finally, we combine the above losses into a single alignment objective to enforce epidemiologically meaningful behavior:
$\mathcal{L}_{\text{align}}=\mathcal{L}_{R_{0}}+\mathcal{L}_{\text{mono}}+\mathcal{L}_{\text{smooth}},\\ \quad\mathcal{L}_{\text{smooth}}=\sum_{k=1}^{K}\sum_{c=1}^{C-2}(\pi_{k,c+2}^{(L)}-2\pi_{k,c+1}^{(L)}+\pi_{k,c}^{(L)})^{2},$ (6)
where $\mathcal{L}_{\text{smooth}}$ is a smoothness regularizer that encourages gradual transitions over time.
3.4 Optimization Scheme
To stably train the model, we employ an alternating optimization strategy. We alternate between freezing the main model to update the prototype embeddings $\mathbf{E}$, and then freezing $\mathbf{E}$ to update the main model parameters. This process is applied across two training phases. During pre-training, we use the self-supervised objectives $\mathcal{L}_{\text{recon}}$ and $\mathcal{L}_{\text{CL}}$, as well as the epidemic-aware regularization $\mathcal{L}_{\text{align}}$, which gives $\mathcal{L}_{\text{pretrain}}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{CL}}+\lambda\mathcal{L}_{\text{align}}$, where $\lambda$ is a hyperparameter. During finetuning on a specific type of pathogen, we adopt the forecasting loss (MSE) in place of the self-supervised objective, giving $\mathcal{L}_{finetune}=\text{MSE}(\mathbf{y}-\mathbf{\hat{y}})+\lambda\mathcal{L}_{\text{align}}$.
@@ 3.4 Optimization Scheme
Table 1: Univariate forecasting results with horizons ranging from 1 to 16 future steps. The lookback window length is set to 36. All models are evaluated over 25 runs, and we report the average MSE and MAE. For CAPE, we also report the 95% confidence interval.
Dataset Horizon Transformer-Based (w/ or w/o pre-train) CAPE
LSTM GRU Dlinear Informer Autoformer MOMENT PatchTST
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Covid 1 32.290 6.749 25.222 5.739 25.799 5.476 28.412 5.778 40.465 6.837 32.026 5.029 26.824 5.159 25.841 (±0.129) 3.611 (±0.011)
2 37.001 7.449 30.856 7.165 27.023 5.873 28.550 4.773 42.969 13.338 31.661 4.123 27.306 5.652 25.413 (±0.076) 3.763 (±0.008)
4 38.129 8.449 29.928 8.065 30.143 6.407 45.663 8.595 46.000 7.960 35.21 5.319 25.756 5.139 24.631 (±0.049) 3.749 (±0.034)
8 45.500 10.680 45.337 10.595 37.733 7.393 55.651 9.945 37.424 9.758 39.633 5.940 38.908 8.998 33.003 (±0.033) 4.827 (±0.005)
16 64.599 12.553 66.860 13.476 55.767 9.259 62.572 14.306 102.196 9.165 51.948 8.116 47.110 7.704 49.838 (±0.050) 7.144 (±0.007)
Avg 43.504 9.176 39.640 9.008 35.293 6.882 44.170 8.679 53.811 9.411 38.096 5.705 33.181 6.530 31.745 (±0.063) 4.619 (±0.014)
ILI USA 1 0.196 0.130 0.221 0.138 0.177 0.123 0.898 0.394 0.805 0.391 0.310 0.169 0.332 0.216 0.174 (±0.003) 0.139 (±0.001)
2 0.281 0.156 0.322 0.167 0.224 0.148 0.395 0.229 0.806 0.399 0.328 0.176 0.283 0.180 0.192 (±0.002) 0.141 (±0.001)
4 0.444 0.197 0.588 0.211 0.305 0.183 0.909 0.400 0.868 0.403 0.434 0.211 0.431 0.257 0.299 (±0.001) 0.171 (±0.000)
8 0.549 0.225 0.771 0.258 0.469 0.234 0.929 0.426 0.899 0.427 0.511 0.222 0.497 0.265 0.469 (±0.001) 0.221 (±0.000)
16 1.515 0.332 0.946 0.287 0.595 0.269 0.690 0.351 0.970 0.410 0.709 0.259 0.651 0.311 0.650 (±0.001) 0.278 (±0.000)
Avg 0.597 0.208 0.569 0.212 0.354 0.191 0.764 0.360 0.870 0.406 0.459 0.207 0.439 0.246 0.357 (±0.001) 0.190 (±0.000)
ILI Japan 1 0.514 0.844 0.552 0.910 0.416 0.880 2.353 1.682 0.715 1.340 0.614 3.459 0.936 1.868 0.328 (±0.001) 0.982 (±0.002)
2 0.758 1.384 0.745 0.904 0.648 0.956 2.446 1.767 0.928 2.786 1.490 4.111 1.531 3.443 0.709 (±0.002) 1.206 (±0.002)
4 1.278 2.733 1.736 3.169 1.253 1.919 2.632 1.884 1.464 4.003 1.542 3.811 1.834 3.355 1.191 (±0.004) 2.029 (±0.002)
8 1.932 1.660 1.948 2.240 1.988 2.196 2.840 1.741 1.925 1.375 2.101 2.314 2.128 1.910 1.792 (±0.002) 1.088 (±0.013)
16 2.118 1.657 2.097 1.550 1.884 1.517 2.490 1.633 2.438 1.799 2.314 1.255 2.265 1.698 1.878 (±0.002) 1.163 (±0.001)
Avg 1.320 1.656 1.415 1.755 1.238 1.493 2.552 1.741 1.494 2.260 1.556 2.99 1.739 2.455 1.179 (±0.002) 1.294 (±0.004)
Measles 1 0.191 1.076 0.202 1.249 0.207 1.022 0.428 2.188 0.699 2.733 0.207 1.270 0.257 1.297 0.111 (±0.005) 0.615 (±0.004)
2 0.230 1.249 0.251 1.147 0.232 1.183 0.479 2.094 0.584 1.672 0.248 1.359 0.418 1.509 0.157 (±0.003) 1.078 (±0.002)
4 0.261 1.153 0.304 1.175 0.297 1.442 1.639 3.510 0.851 2.005 0.296 1.395 0.459 1.512 0.188 (±0.003) 1.352 (±0.001)
8 0.415 2.007 0.392 1.703 0.468 1.938 0.592 2.627 1.171 2.767 0.476 1.771 0.721 2.558 0.406 (±0.002) 2.002 (±0.000)
16 0.696 2.431 0.729 2.695 0.953 2.864 2.098 3.793 1.922 3.924 0.763 2.747 1.271 3.164 0.883 (±0.001) 2.836 (±0.000)
Avg 0.358 1.583 0.375 1.594 0.431 1.690 1.047 2.843 1.046 2.620 0.398 1.708 0.625 2.008 0.349 (±0.003) 1.576 (±0.002)
Dengue 1 0.583 1.579 0.627 1.343 0.503 1.074 0.627 1.697 1.556 2.456 0.630 1.435 2.056 2.600 0.367 (±0.011) 1.282 (±0.003)
2 0.634 1.417 0.676 1.604 0.566 1.337 0.905 2.111 1.827 2.622 0.690 1.597 0.925 1.567 0.317 (±0.006) 1.272 (±0.006)
4 0.823 2.137 0.984 1.940 0.845 1.767 1.170 2.184 2.546 2.887 0.938 1.827 1.419 2.226 0.508 (±0.003) 1.534 (±0.003)
8 1.534 2.758 1.375 2.623 1.488 2.410 1.392 2.428 3.679 3.322 1.504 2.283 1.581 2.261 1.169 (±0.004) 2.104 (±0.002)
16 2.561 3.078 2.745 2.914 2.861 3.003 3.841 3.454 4.734 3.581 2.768 2.932 4.923 3.618 2.512 (±0.003) 2.784 (±0.006)
Avg 1.227 2.194 1.281 2.085 1.252 1.918 1.587 2.375 2.868 2.973 1.306 2.015 2.181 2.454 0.975 (±0.005) 1.795 (±0.004)